OpenStudy (anonymous):

Which value of k makes (2x + 3)(kx – 5) = 2kx^2 – x – 15 true? A. –3 B. –1 C. 1 D. 3

OpenStudy (anonymous):

Can anybody help?

OpenStudy (anonymous):

well you're going to have to expand the left side on this one

OpenStudy (anonymous):

then see if the x's will cancel somehow so you can solve for k

OpenStudy (anonymous):

I have to try the values right?

OpenStudy (anonymous):

well you could do that, but it is not necessary

OpenStudy (anonymous):

then how should I solve it?:(

OpenStudy (shubhamsrg):

Simplify LHS, tell us what you get

OpenStudy (anonymous):

2kx^2 +3kx - 10x -15 = 2kx^2 - x - 15 k(2x^2 + 3x) - 10x - 15 = 2kx^2 - x = 15 subtract 2kx^2 - 10x - 15 from both sides 3kx = 9x divide by 3x on both sides k = 3

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

you're welcome, that one was a little tricky, just know that you have to be able to get rid of the x variable somehow and leave a constant behind